# Distributed Maximum Flow in Planar Graphs

### v1

@@ Abstract
Given a directed planar graph $G$ with positive and negative edge-lengths and hop-diameter $D$, our key result is an $\tilde{O}(D^{2})$-round algorithm [^1] The $\tilde{O}(\cdot)$ notation is used to omit $\text{poly}\log n$ factors. for Single Source Shortest Paths on $G^{*}$, which then implies an $\tilde{O}(D^{2})$-round algorithm for Maximum $st$-Flow on $G$. Prior to our work, no $\tilde{O}(\text{poly}(D))$-round algorithm was known for Maximum $st$-Flow. We further obtain a $D\cdot n^{o(1)}$-rounds $(1-\epsilon)$-approximation algorithm for Maximum $st$-Flow on $G$ when $G$ is undirected and $s$ and $t$ lie on the same face. Finally, we give a near optimal $\tilde{O}(D)$-round algorithm for computing the weighted girth of $G$. We believe that the toolkit developed in this paper for exploiting planar duality will be used in future distributed algorithms for various other classical problems on planar graphs (as happened in the centralized setting).
@@ 1 Introduction
In directed planar graphs with integral edge-capacities, it is known from the 80’s [44] that the maximum $st$-flow can be found by solving $\log\lambda$ instances of Single Source Shortest Paths (SSSP) with positive and negative edge-lengths on the dual graph $G^{*}$, where $\lambda$ is the maximum $st$-flow value. We answer 1.1 in the affirmative by designing a $\tilde{O}(D^{2})$-round SSSP algorithm on the dual graph $G^{*}$. Our algorithm works in the most general setting w.h.p. [^2] W.h.p. stands for a probability of $1-1/{n^{c}}$ for an arbitrary fixed constant $c>0$. (i.e. when $G^{*}$ is directed and has positive and negative integral edge-lengths) and matches the fastest known exact SSSP algorithm in the primal graph. We show:
@@
Minimum $st$-cut. By the well-known Max Flow Min Cut theorem of [7], our flow algorithms immediately give the value (or approximate value) of the minimum $st$-cut. We show that they can be extended to compute a corresponding bisection and the cut edges without any overhead in the round complexity. Moreover, since our exact flow algorithm works with directed planar graph, it admits a solution to the directed minimum $st$-cut problem. To the best of our knowledge, prior to our work, $\tilde{O}(\text{poly}(D))$-round CONGEST algorithms for the minimum $st$-cut problem were known only for general graphs with constant cut values by [33]. This is in contrast to the global minimum cut problem that can be solved in $\tilde{O}(D)$ rounds in undirected planar graphs [14, 18].
Primal weighted girth via dual cuts. A distance parameter of considerable interest is the network girth. For unweighted graphs, the girth is the length of the smallest cycle in the graph. For weighted graphs, the girth is the cycle of minimal total edge weight. Distributed girth computation has been studied over the years mainly for general $n$-vertex unweighted graphs. Frischknecht, Holzer and Wattenhofer [8] provided an $\Omega(\sqrt{n})$-round lower bound for computing a $(2-\epsilon)$ approximation of the unweighted girth. The state-of-the-art upper bound for the unweighted girth problem is a $(2-\epsilon)$ approximation in $\tilde{O}(n^{2/3}+D)$ rounds, obtained by combining the works of Peleg, Roditty and Tal [36] and Holzer and Wattenhofer [21]. The weighted girth problem has been shown to admit a near-optimal lower bound of $\tilde{\Omega}(n)$ rounds in general graphs [22, 28]. Turning to planar graphs, Parter [34] devised a $\tilde{O}(D^{2})$ round algorithm for computing the weighted girth in directed planar graphs via SSSP computations. For undirected and unweighted planar graphs, the (unweighted) girth can be computed in $\tilde{O}(D)$ rounds by replacing the $\tilde{O}(D^{2})$-round SSSP algorithm by a $O(D)$-round BFS algorithm. In light of this gap, we ask:
@@
There is a randomized distributed algorithm that given an $n$-vertex undirected weighted planar network $G$ with hop-diameter $D$, computes the weighted girth (and finds a corresponding cycle) w.h.p. in $\tilde{O}(D)$ rounds.
@@ 2.2 SSSP in the Dual via Distance Labels
Bounded diameter decomposition (BDD). Is a distributed hierarchical decomposition for planar graphs, which plays an analogous role to the centralized recursive separator decomposition, which was devised by Li and Parter [26] by carefully using and extending a distributed planar separator algorithm of Ghaffari and Parter [17]. The BDD has some useful properties that are specific to the distributed setting. In particular, all the separators and subgraphs obtained in each recursive level have small diameter of $\tilde{O}(D)$ and are nearly disjoint, which allows to broadcast information in all of them efficiently in $\mathsf{CONGEST}$. An immediate application of the BDD is a distributed distance labeling algorithm for primal planar graphs that follows the intuition above. In addition, BDDs have found other applications in $\mathsf{CONGEST}$ algorithms on the primal graph $G$ (e.g. diameter approximation, routing schemes and reachability [6, 26, 34]).
@@ 2.3 Flow Assignments and Cuts
Roadmap.
The rest of the paper is organized as follows. Section 3 discusses preliminaries. Section 4 gives an overview of our minor-aggregation simulation on the dual graph and its applications for the weighted girth, where full details and proofs appear in Appendix B. Our labeling scheme for the dual graph appears in Section 5, and our applications for max $st$-flow and min $st$-cut appear in Section 6. See also Figure 6.
Dual Minor-Aggregation (Section 4, Appendix B)Undirected Weighted Girth (Theorem 1.5)Approximate Dual SSSPExtended BDD (Section 5.1)Dual Distance Labeling (Section 5.2)Exact Dual SSSP(Section 5.3)Primal Max $st$-Flow(Section 6.1)Primal Min $st$-Cut(Section 6.2)
@@ 3 Preliminaries
Fact 3.1 (Cycle-Cut Duality).
A set of edges $C$ is a simple cycle in a connected planar graph $G$ if and only if $C$ is a simple cut in $G^{*}$ (a cut $C$ is said to be simple if when removed, the resulting graph has exactly two connected components).
@@ 7. Distributed knowledge:
    Each node $f$ of $X^{*}$ has an $\tilde{O}(1)$ ID, s.t. the vertices $v$ of $X$ lying on $f$ know its ID. Moreover, $v$ knows whether $f$ corresponds to a face, a face-part or the critical face of $G$.
@@ 5.1.1 The Few Face-parts Property
We start by proving Property 9. Mainly, we show,
@@ 5.1.3 Distributed Knowledge Properties
We aim to assign each face (face-part) in any bag $X$ a unique $\tilde{O}(1)$ identifier and learn for each vertex $v\in X$, the set of faces (face-parts) containing it and its adjacent edge copies. This is done by keeping track of $G$’s edge copies along the decomposition. We start with (1) a local procedure that allows the endpoints of each edge copy $e$ to learn all the bags $X$ that contain $e$, then, (2) the endpoints of each edge copy $e$ learn the face of $G$ that $e$ participates in, and (3) we show how to extend this for the next level of $\mathcal{T}$. For a proof see Appendix C. This allows us to (distributively) learn the nodes of $X^{*}$, after which, we have the necessary information to learn the edges of $X^{*}$. Having:
@@
We get the following lemma and corollary proving property 13 of the BDD:
@@
There is a $\tilde{O}(D)$-round distributed algorithm that assigns unique $\tilde{O}(1)$-bit IDs to all faces and face-parts in all bags $X\in\mathcal{T}$. By the end of the algorithm, each vertex $v\in G$ learns a list (of IDs) of the faces and face-parts it lies on for each bag $X$ that contains it. Moreover, $v$ learns for each incident copy of an edge $e^{\prime}$ the (ID of) the face or face-part that $e^{\prime}$ participates in for each bag $X$ that contains $e^{\prime}$.
@@ 6 Primal Maximum st-Flow and Minimum st-Cut
In this section we show how to use the labeling algorithm of Section 5 to compute the maximum $st$-flow in $G$, and how to use the minor-aggregation model simulation on $G^{*}$ of Section 4 to approximate the maximum $st$-flow in $G$ in the special case when $G$ is undirected and both $s$ and $t$ lie on the same face in the given planar embedding. I.e, when $G$ is $st$-planar.
@@ 6.1 Maximum st-Flow
It remains only to show that we can apply [39]’s algorithm (Theorem 5.1, Algorithms 3 and 4 in their paper), in order to obtain smooth $(1+\epsilon)$ approximate single source shortest paths from $f_{1}$ in $G^{*}_{virt}=(G\cup\{e\})^{*}$ by invoking the algorithm of [41] that we use for approximating distances, which we refer to as the oracle. We note that [39]’s algorithm assumes real weights in $[1,n^{O(1)}]$, whereas we work with integral weights, however, they suggest a fix (by scaling) for this that appears in Appendix C of their paper. Note, scaling can be implemented efficiently in one minor-aggregation round, henceforth, we consider this issue resolved.
@@ 6.2 Minimum st-Cut
Theorem 6.1 (Distributed Exact Directed Minimum $st$-Cut in Planar Graphs).
@@
Theorem 6.2 (Distributed Approximate Undirected Minimum $st$-Cut in $st$-Planar Graphs).

### v3

@@ Abstract
Given a directed planar graph $G$ with positive and negative edge-lengths and hop-diameter $D$, our key result is an $\tilde{O}(D^{2})$-round algorithm [^1] The $\tilde{O}(\cdot)$ notation is used to omit $\text{poly}\log n$ factors. for Single Source Shortest Paths on $G^{*}$. This algorithm implies an $\tilde{O}(D^{2})$-round algorithm for Maximum $st$-Flow in $G$, and an $\tilde{O}(D^{2})$-round algorithm for directed Global Minimum Cut in $G$. Prior to our work, no $\tilde{O}(\text{poly}(D))$-round algorithms were known for these problems. When $G$ is undirected, we further obtain a near optimal $\tilde{O}(D)$-round algorithm for computing the weighted girth of $G$, and a $D\cdot n^{o(1)}$-rounds $(1-\epsilon)$-approximation algorithm for Maximum $st$-Flow in $G$ when $s$ and $t$ lie on the same face.
@@
We believe that the toolkit developed in this paper for exploiting planar duality will be used in future distributed algorithms for various other classical problems on planar graphs (as happened in the centralized setting).
@@ 1 Introduction
In directed planar graphs with integral edge-capacities, it is known from the 80’s [45] that the maximum $st$-flow can be found by solving $\log\lambda$ instances of Single Source Shortest Paths (SSSP) with positive and negative edge-lengths on the dual graph $G^{*}$, where $\lambda$ is the maximum $st$-flow value. We answer 1.1 in the affirmative by designing a $\tilde{O}(D^{2})$-round SSSP algorithm on the dual graph $G^{*}$. Our algorithm works w.h.p. [^2] W.h.p. stands for a probability of $1-1/{n^{c}}$ for an arbitrary fixed constant $c>0$., it works in the most general setting (i.e. when $G^{*}$ is directed and has positive and negative integral edge-lengths) and it matches the fastest known exact SSSP algorithm in the primal graph. We show:
@@
Minimum $st$-cut. By the well-known Max Flow Min Cut theorem of [7], our flow algorithms immediately give the value (or approximate value) of the minimum $st$-cut. We show that they can be extended to compute a corresponding bisection and the cut edges without any overhead in the round complexity. Moreover, since our exact flow algorithm works with directed planar graphs, it admits a solution to the directed minimum $st$-cut problem. To the best of our knowledge, prior to our work, $\tilde{O}(\text{poly}(D))$-round CONGEST algorithms for the minimum $st$-cut problem were known only for general graphs with constant cut values by [34]. This is in contrast to the global (undirected) minimum cut problem that can be solved in $\tilde{O}(D)$ rounds in planar graphs [14, 18].
Directed global minimum cut. The directed global minimum weight cut problem is given a weighted directed graph and asks to find a bisection $(S,V\setminus S)$ of the vertex set, such that, the set of edges leaving $S$ to $V\setminus S$ is of minimal total weight. To our knowledge, this problem was not studied in the $\mathsf{CONGEST}$ model and is open even for approximations, as opposed to the (directed or undirected) minimum $st$-cut problem [4, 34] and the undirected global minimum cut problem [18, 5, 14]. In light of efficient planar centralized results [31, 46] and motivated by the importance of the problem, we ask:
@@
Is it possible to compute the directed global minimum cut in planar graphs within $\tilde{O}(\text{poly}(D))$ rounds?
We answer 1.4 in the affirmative by showing an extension to our SSSP algorithm in the dual, allowing us to compute the minimum weight directed cycle in the dual. By duality, the same set of edges corresponds to the desired cut and determines a corresponding bisection in the primal, all within the same $\tilde{O}(D^{2})$ bound on the round complexity:
Theorem 1.5 (Planar Directed Global Minimum Cut).
Given a directed planar graph $G$ with non-negative integral edge weights, there is an $\tilde{O}(D^{2})$ round randomized algorithm that finds a global directed minimum cut w.h.p..
Primal weighted girth via dual cuts. A distance parameter of considerable interest is the network girth. For unweighted graphs, the girth is the length of the shortest cycle in the graph. For weighted graphs, the girth is the cycle of minimal total edge weight. Distributed girth computation has been studied over the years mainly for general $n$-vertex unweighted graphs. Frischknecht, Holzer and Wattenhofer [8] provided an $\Omega(\sqrt{n})$-round lower bound for computing a $(2-\epsilon)$ approximation of the unweighted girth. The state-of-the-art upper bound for the unweighted girth problem is a $(2-\epsilon)$ approximation in $\tilde{O}(n^{2/3}+D)$ rounds, obtained by combining the works of Peleg, Roditty and Tal [37] and Holzer and Wattenhofer [21]. The weighted girth problem has been shown to admit a near-optimal lower bound of $\tilde{\Omega}(n)$ rounds in general graphs [22, 28]. Turning to planar graphs, Parter [35] devised a $\tilde{O}(D^{2})$ round algorithm for computing the weighted girth in directed planar graphs via SSSP computations. For undirected and unweighted planar graphs, the (unweighted) girth can be computed in $\tilde{O}(D)$ rounds by replacing the $\tilde{O}(D^{2})$-round SSSP algorithm by a $O(D)$-round BFS algorithm. In light of this gap, we ask:
Question 1.6.
@@
There is a randomized distributed algorithm that given an $n$-vertex undirected weighted planar network $G$ with hop-diameter $D$, computes the weighted girth and finds the edges of a shortest cycle w.h.p. in $\tilde{O}(D)$ rounds.
@@ 2.2 SSSP in the Dual via Distance Labels
Bounded diameter decomposition (BDD). The BDD is a distributed hierarchical decomposition for planar graphs, which plays an analogous role to the centralized recursive separator decomposition. It was devised by Li and Parter [26] by carefully using and extending a distributed planar separator algorithm of Ghaffari and Parter [17]. The BDD has some useful properties that are specific to the distributed setting. In particular, all the separators and subgraphs obtained in each recursive level have small diameter of $\tilde{O}(D)$ and are nearly disjoint, which allows to broadcast information in all of them efficiently in $\mathsf{CONGEST}$. An immediate application of the BDD is a distributed distance labeling algorithm for primal planar graphs that follows the intuition above. In addition, BDDs have found other applications in $\mathsf{CONGEST}$ algorithms on the primal graph $G$ (e.g. diameter approximation, routing schemes and reachability [6, 26, 35]).
@@ 2.3 Flow Assignments and Cuts
Directed global minimum cut. We exploit our techniques to find also the directed global minimum cut. The general idea is simple, we use the cycle-cut duality: finding the (directed) minimum cut in $G$ amounts to finding the minimum weight (directed) simple cycle $C$ in $G^{*}$. Thus, given an extended BDD of $G$ and having distance labels already computed in $\tilde{O}(D^{2})$ rounds for all dual bags in the BDD, we have that any shortest cycle of $G^{*}$ either intersects a dual separator node or is entirely contained in one of $G^{*}$’s child bags. Thus, in order to find $C$, we take the minimum over all: (i) directed cycles that intersect separator nodes of $G^{*}$ (deduced by pairwise decoding labels of the separator nodes), and (ii) minimum directed cycles in $G^{*}$’s child bags (found recursively). However, we note that the duality does not hold in general directed planar graphs, which complicates the above approach. I.e., in order for the duality to hold, we have to work with a graph where for each edge, its reversal dart (an edge in the opposite direction) is present, so we add those darts. Then, we should ensure that the cycles we consider are simple darts cycles (i.e. a cycle where no edge and its reversal dart appear together), which we do by investigating two options related to the dual separator. Finding the cut bisection is done simply by detecting connected components (in the undirected sense) after the removal of the cut edges and their reversal darts.
2.4 Organization, Summary and Open Questions
The rest of the paper is organized as follows. Section 3 discusses preliminaries. Section 4 gives an overview of our minor-aggregation simulation on the dual graph and its applications for the weighted girth, where full details and proofs appear in Appendix B. Our labeling scheme for the dual graph appears in Section 5, and our applications for max $st$-flow, min $st$-cut appear in Section 6, while the directed global min-cut appears in Section 7. See also Figure 6.
Dual Minor-Aggregation (Section 4, Appendix B)Undirected Weighted Girth (Theorem 1.7)Approximate Dual SSSPExtended BDD (Section 5.1)Dual Distance Labeling (Section 5.2)Exact Dual SSSP(Section 5.3)Primal Max $st$-Flow(Section 6.1)Primal Directed Global Min-Cut(Section 7)Primal Min $st$-Cut(Section 6.2)
@@ 2.4 Organization, Summary and Open Questions
Next, we summarize the techniques by their advantages and limitations, then, we conclude with an interesting open question.
Minor-Aggregation.
In general, the minor aggregation technique is useful when dealing with undirected graphs and especially when we want to simulate some related graph to the graph of input, which includes a small number (usually $\tilde{O}(1)$) virtual nodes. I.e., nodes that all vertices of $G$ simulate. Albeit the dual graph $G^{*}$ is entirely virtual, we extend previous work [17] and show that the round complexity of simulating a minor aggregation round on it matches the near-optimal bound [41] known for $G$, which allows then to add up to $\tilde{O}(1)$ virtual nodes to the dual graph that are not necessarily related to the faces of $G$. Then, applying black-box algorithms on $G^{*}$ after manipulating its topology (adding/ replacing nodes with virtual nodes), constitutes a distributed implementation of classic centralized reductions, some of which are immediate (e.g. min-cut in $G^{*}$ gives the girth of $G$) and some of which require more work (e.g. shortest paths with positive edge-lengths in $G^{*}$ implies maximum $st$-flow in $G$ when $s,t$ lie on the same face).
This closes the problem of minor-aggregation on the dual and expands the algorithmic design toolkit to include all the very useful efficient minor-aggregation procedures (e.g. subtree sums, Heavy-Light decompositions [18], etc) and even complicated algorithms (e.g. min-cut [18], approximate undirected shortest path [41]) for the dual.
Recursive Decompositions.
Our recursive decomposition based on the BDD of [26] provides a framework to divide-and-conquer in $G^{*}$ making this approach feasible albeit the difficulties encountered. In this scenario, usually, one solves the problem in a brute-force manner on the leaf-bags then broadcasts some information to the parent bags in order to assemble a solution for the parent bag, and so on, bubbling up to the root of the decomposition $G^{*}$. The most obvious example is the SSSP algorithm (Section 5) we give to compute maximum $st$-flow in $G$. An additional example is given in Section 7, where we show how to use the distance labels computed in $G^{*}$ in a non-trivial manner to compute its minimal weight (directed) cycle, which results with the global directed minimum cut of $G$.
Having a recursive separator decomposition for $G^{*}$, opens the question also to other kinds of graph decompositions known to the primal graph with efficient distributed constructions (e.g. shortest-path separator recursive decomposition [26]). Although the dual decomposition itself can be constructed and distributively learned in near-optimal time, the barrier for a better maximum flow algorithm lies within the dual SSSP algorithm, which is discussed next.
Open Questions.
Since the round complexity of the algorithm that we devise for directed single source shortest paths with positive and negative edge-lengths for $G^{*}$ matches the state-of-the-art for $G$, then, perhaps the most obvious question is to obtain better SSSP algorithms for the primal graph $G$ (where we do not have the challenges of simulating an entirely virtual graph as $G^{*}$). I.e., in $G$, one should probably first improve the $\tilde{O}(D^{2})$ bound of the (easier) directed SSSP problem (even with non-negative weights) before trying to achieve such an improvement for (the harder) maximum $st$-flow problem.
It seems that this is the bottleneck to overcome in order to obtain better distributed planar algorithms for various optimization problems, since most such problems in $G$ reduce to computing distances either in $G$ or in $G^{*}$.
It is intuitive to believe that an improvement upon the bound of $\tilde{O}(D^{2})$ is possible, because every step in the SSSP algorithm is near-optimal except when we simply broadcast the labels of all separator vertices ($\tilde{\Theta}(D^{2})$ bits), if one can compress those distances to $\tilde{o}(D^{2})$ bits or manage to compute SSSP in a better round complexity via a different approach, then likely better algorithms for many planar optimization problems would follow. An indication that this is possible was given by Parter [35], where she shows near-optimal algorithms for planar reachability problems using the BDD, by compressing the reachability labels of all separator vertices into $\tilde{O}(D)$ bits before the broadcast step.
@@ 3 Preliminaries
Fact 3.1 (Undirected Cycle-Cut Duality [25]).
A set of edges $C$ is a simple cycle in an embedded connected planar graph $G$ if and only if $C$ is a simple cut in $G^{*}$ (a cut $C$ is said to be simple if when removed, the resulting graph has exactly two connected components).
@@ 7. Distributed knowledge:
    Each node $f$ of $X^{*}$ has an $\tilde{O}(1)$ ID, s.t. the vertices $v$ of $X$ lying on $f$ know its ID. Moreover, $v$ knows whether $f$ corresponds to a face, a face-part or the critical face of $X$. The IDs are such that the ID of a face (face-part) $f$ of $X$ is contained in the ID of face-parts $f^{\prime}$ resulting from partitioning $f$ in child bags of $X$.
@@ 5.1.1 The Few Face-parts Property
We start by proving Property 9.
@@ 5.1.3 Distributed Knowledge Properties
We aim to assign each face (face-part) in any bag $X$ a unique $\tilde{O}(1)$ identifier and learn for each vertex $v\in X$, the set of faces (face-parts) containing it and its adjacent edge copies. This is done by keeping track of $G$’s edge copies along the decomposition. We start with (1) a local procedure that allows the endpoints of each edge copy $e$ to learn all the bags $X$ that contain $e$, then, (2) the endpoints of each edge copy $e$ learn the face of $G$ that $e$ participates in, and (3) we show how to extend this for the next level of $\mathcal{T}$. For a proof see Appendix C. This allows us to (distributively) learn the nodes of $X^{*}$, after which, we have the necessary information to learn the edges of $X^{*}$.
@@
The IDs assigned to faces and face parts are defined such that the ID of a face (face-part) $f$ of $X$ is contained in the ID of face-parts resulting from partitioning $f$ in child bags of $X$. Thus, we get the following lemma and corollary proving property 13 of the BDD:
@@
There is a $\tilde{O}(D)$-round distributed algorithm that assigns unique $\tilde{O}(1)$-bit IDs to all faces and face-parts in all bags $X\in\mathcal{T}$. By the end of the algorithm, each vertex $v\in G$ learns a list (of IDs) of the faces and face-parts it lies on for each bag $X$ that contains it. Moreover, $v$ learns for each incident copy of an edge $e^{\prime}$ the (ID of) the face or face-part that $e^{\prime}$ participates in for each bag $X$ that contains $e^{\prime}$. The ID of a face (face-part) $f$ of $X$ is contained in the ID of face-parts $f^{\prime}$ resulting from partitioning $f$ in child bags of $X$.
@@ 6 Primal Maximum st-Flow and Minimum st-Cut
In this section we show how to use the labeling algorithm of Section 5 to compute the maximum $st$-flow in $G$, in addition, how to use the minor-aggregation model simulation on $G^{*}$ of Section 4 to approximate the maximum $st$-flow in $G$ in the special case when $G$ is undirected and both $s$ and $t$ lie on the same face in the given planar embedding. I.e, when $G$ is $st$-planar. The maximum $st$-flow results naturally extend to minimum $st$-cut algorithms.
@@ 6.1 Maximum st-Flow
that we can apply [40]’s algorithm (Theorem 5.1, Algorithms 3 and 4 in their paper), in order to obtain smooth $(1+\epsilon)$ approximate single source shortest paths from $f_{1}$ in $G^{*}_{virt}=(G\cup\{e\})^{*}$ by invoking the algorithm of [42] that we use for approximating distances, which we refer to as the oracle. We note that [40]’s algorithm assumes real weights in $[1,n^{O(1)}]$, whereas we work with integral weights, however, they suggest a fix (by scaling) for this that appears in Appendix C of their paper. Note, scaling can be implemented efficiently in one minor-aggregation round, henceforth, we consider this issue resolved.
@@ 6.2 Minimum st-Cut
Theorem 6.1 (Distributed Exact Planar Directed Minimum $st$-Cut).
@@
Theorem 6.2 (Distributed Approximate $st$-Planar Undirected Minimum $st$-Cut).